OpenStudy (anonymous):

Find the product of (x + 3)(x - 3). x2 - 6x + 9 x2 + 6x + 9 x2 + 9 x2 - 9

OpenStudy (anonymous):

Same process.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

im thinking its a

OpenStudy (anonymous):

x*x=x^2 Correct?

OpenStudy (anonymous):

x times x = x^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3*3=9

OpenStudy (anonymous):

3+3=6

OpenStudy (anonymous):

Therefore, your answer will be.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@AvatarKorra, D, isn't the right answer.

OpenStudy (anonymous):

@Alexandra675 correct!

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

???? isnt D correct

OpenStudy (anonymous):

Got the addition and subtraction sign all mixed up lol

OpenStudy (anonymous):

what did u come up with vzfreakz

OpenStudy (anonymous):

It was wrong...

OpenStudy (anonymous):

What? How was it wrong?

OpenStudy (anonymous):

I simplified this. Well, do you have another attempt on this question?

OpenStudy (anonymous):

no not really but I don't get how it was wrong and the trinomial question was wrong too...

OpenStudy (anonymous):

It doesn't say the answer either...

OpenStudy (anonymous):

it is the theorem u asked earlier Alexandra675 it is always a binomial

OpenStudy (anonymous):

I simplified this correctly. Not quite sure what's wrong.

OpenStudy (anonymous):

B is wrong vzfreakz

OpenStudy (anonymous):

@AvatarKorra yes, I think I figured that out already Avatar..

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

@AvatarKorra will explain it to you.

OpenStudy (anonymous):

A? Please don't tell me its A

OpenStudy (anonymous):

How many chances you got?

OpenStudy (anonymous):

thank u vzfreakz for at last giving me the time. (a+b)(a-b) = asquared -bsquared

OpenStudy (anonymous):

Its D. cauz x times x equals xsquared and then 3.3 gives 9

OpenStudy (anonymous):

so it is xsquared - 9 try that

OpenStudy (anonymous):

*dies in a hole*

OpenStudy (anonymous):

?

OpenStudy (anonymous):

was this ans also incorrect???

OpenStudy (anonymous):

no I have no more tries but I cant belive it lol

OpenStudy (anonymous):

ok